# MFCPL: Multimodal Federated Cross Prototype Learning for Missing Modalities and Data Heterogeneity

### v1

@@ 3.2 Multimodal Prototype Federated Learning
Prior research works in federated prototype learning Tan et al. (2022b); Mu et al. (2023) have focused on scenarios where clients possess unimodal data. In the context of MFL under missing modalities, multiple types of clients exist: those missing modality $A$, those missing modality $B$, and those with complete modalities $A$ and $B$. Simply aggregating unimodal prototypes based on modality may introduce a bias toward the dominant modality and cause the performance degradation in FL under severely missing modalities. This prompts us to leverage prototypes from all modalities to develop a generalized model for different types of clients.
Complete Prototypes.
@@
$\{r_{A}=g_{1}(e_{A}),r_{B}=g_{1}(e_{B}),r_{AB}=g_{1}(e_{AB})\}\in\mathbb{R}^{d}.$ (2)
We define the $k^{th}$ class prototype from different types of clients as:
$p^{k}_{Ai}=\frac{1}{|S^{k}_{Ai}|}\sum_{i\in S^{k}_{Ai}}r_{Ai},$ (3)
$p^{k}_{Bi}=\frac{1}{|S^{k}_{Bi}|}\sum_{i\in S^{k}_{Bi}}r_{Bi},$ (4)
$p^{k}_{ABi}=\frac{1}{|S^{k}_{ABi}|}\sum_{i\in S^{k}_{ABi}}r_{ABi},$ (5)
where $S^{k}_{Ai}$, $S^{k}_{Bi}$ and $S^{k}_{ABi}$ are subsets of $D_{i}$ belonging to $k^{th}$ class of different client types, respectively. As complementary to unimodal prototypes, we initially introduce complete prototypes by averaging local prototypes from different types of clients in the shared representation space $\mathbb{R}^{d}$. Specifically, we formulate the complete prototypes on the server as follows:
$P^{k}=\frac{1}{N}\sum_{i\in N}\{p^{k}_{Ai},p^{k}_{Bi},p^{k}_{ABi}\}\in\mathbb{R}^{d},$ (6)
where $P^{k}$ denotes the complete prototypes belonging to class $k$. Specifically, our complete prototypes guide the local training process at two levels: modality-specific and modality-shared representations.
3.3 Multimodal Federated Cross Prototype Learning
Cross-Modal Prototypes Regularization.
Similar to other traditional FL methods, MFL encounters the challenge of data heterogeneity, and further complicated by the issue of missing modalities. In response to these challenges, we introduce a regularization loss that narrows the gap between local and global models, simultaneously providing diverse modality knowledge for clients with missing modalities at the modality-shared representation level. Drawing inspiration from FedProto Tan et al. (2022b), we incorporate a regularization loss between local projected multimodal representations $r_{i}$ and complete prototypes $P^{k}$ within the same semantic class $k$. However, our approach differs from FedProto, as our regularization loss serves as an extension to address the challenges posed by missing modalities in the context of multimodal federated learning. To implement this regularization, we utilize $\ell_{2}$ distance and introduce the Cross-Modal Prototypes Regularization (CMPR) as follows:
$\mathcal{L}_{CMPR}=\sum_{k}\|r^{k}_{i}-P^{k}\|_{2}^{2}$ (7)
where $r^{k}_{i}$ is the local projected multimodal representations of $k$ semantic class.
@@ 3.3 Multimodal Federated Cross Prototype Learning
3:        $\theta^{i}_{t},p_{Ai},p_{Bi},p_{ABi}\leftarrow\textbf{LocalUpdate}(\theta_{t},P)$
@@
9:     for each batch $\in D_{i}=\{{x}^{A}_{i},{x}^{B}_{i},{y}_{i}\}$ do
10:        $z_{Ai}=f_{A}(x^{A}_{i}),z_{Bi}=f_{B}(x^{B}_{i})$
11:        $e_{i}=h_{\text{self-attention}}(z_{Ai},z_{Bi}),r_{i}=g_{1}(e_{i})$
12:        $z^{\prime}_{Ai}=g_{2}(z_{Ai}),z^{\prime}_{Bi}=g_{2}(z_{Bi})$
13:        $\mathcal{L}_{CMPR}\leftarrow(r_{i},P)$ in Eq. 7
14:        $\mathcal{L}_{CMPC}\leftarrow(z^{\prime}_{Ai},z^{\prime}_{Bi},P)$ in Eq. 11
15:        $\mathcal{L}_{CMA}\leftarrow(z^{\prime}_{Ai},z^{\prime}_{Bi})$ in Eq. 12
@@
20:  $p_{Ai}=[p^{1}_{Ai},p^{2}_{Ai},\dots,p^{k}_{Ai}]$
21:  $p_{Bi}=[p^{1}_{Bi},p^{2}_{Bi},\dots,p^{k}_{Bi}]$
22:  $p_{ABi}=[p^{1}_{ABi},p^{2}_{ABi},\dots,p^{k}_{ABi}]$
23:  return $\theta^{i}_{t},p_{Ai},p_{Bi},p_{ABi}$
Cross-Modal Prototypes Contrastive.
@@
We envision that modality-specific representations in clients with missing modalities can capture rich and diverse modality knowledge from the complete prototype. The modality-specific features of local data instances should align closely with their respective complete prototypes within the same semantic class while exhibiting dissimilarity to complete prototypes associated with different semantic classes. Regarding the data instances $\{{x}^{A}_{i},{x}^{B}_{i},{y}_{i}\}$, we initially encode them using encoder $f_{m}$ to acquire the modality-specific feature vectors $z_{Ai}=f_{A}({x}^{A}_{i}),z_{Bi}=f_{B}({x}^{B}_{i})\in\mathbb{R}^{n}$. To facilitate alignment between the modality-specific features and complete prototypes, we employ the projection head $g_{2}$ to project the modality-specific feature vectors of each modality into the shared representation space:
$z^{\prime}_{Ai}=g_{2}(z_{Ai}),z^{\prime}_{Bi}=g_{2}(z_{Bi})\in\mathbb{R}^{d},$ (8)
@@
Let $p$ be the corresponding complete prototype $p\in P$, $C^{k}=P-P^{k}$ represents the prototypes with different semantic class from the local instances. Subsequently, the Cross-Modal Prototypes Contrastive (CMPR) loss where all modalities are available is expressed as follows:
$\mathcal{L}_{ConA}=-\log\frac{\sum_{p\in P^{k}}\exp(s(z^{\prime}_{Ai},p)/\tau)}{\sum_{p\in P}\exp(s(z^{\prime}_{Ai},p)/\tau)}$ (9)
$\mathcal{L}_{ConB}=-\log\frac{\sum_{p\in P^{k}}\exp(s(z^{\prime}_{Bi},p)/\tau)}{\sum_{p\in P}\exp(s(z^{\prime}_{Bi},p)/\tau)}$ (10)
$\mathcal{L}_{CMPC}=\mathcal{L}_{ConA}+\mathcal{L}_{ConB}$ (11)
@@
where $\mathrm{s}(u,v)=u^{\top}v/\lVert u\rVert\lVert v\rVert$ denotes the cosine similarity between the projected modality-specific feature and complete prototypes, and $\tau$ is the contrastive hyper-parameter. In instances where clients lack modality A or B, the contrastive loss with $z^{\prime}_{Ai}$ or $z^{\prime}_{Bi}$ will be excluded. The goal of Eq. (11) is to encourage the modality-specific representations from clients with missing modalities to closely align with the corresponding complete prototypes in the shared embedding space, thereby enhancing the generalization of these clients. In this component, we utilize the projection head to facilitate alignment between feature vectors at different levels: the complete prototype and modality-specific representations. The effectiveness of the projection head is further discussed in the Experiment subsection 4.4.
Cross-Modal Alignment.
@@
While the complete prototypes contribute diverse modality knowledge to address data heterogeneity and missing modalities at modality-specific and modality-shared representation levels, simply imputing missing data with zeros for the absent modality can introduce noise. We introduce an alignment loss between modality-specific representations from different modalities to alleviate noise arising from representations of missing modalities. We apply the alignment loss to the projected modality-specific representations. To achieve this goal, we utilize the $\ell_{2}$ distance and formulate the distance loss on two modality-specific representations as:
$\mathcal{L}_{CMA}=\sum_{i\in D_{i}}\|z^{\prime}_{Ai}-z^{\prime}_{Bi}\|_{2}^{2},$ (12)
@@
where $z^{\prime}_{Ai}=g_{2}(z_{Ai})$, $z^{\prime}_{Bi}=g_{2}(z_{Bi})\in\mathbb{R}^{d}$ represent the projected modality-specific representations for modalities A and B, respectively. Our objective is to align the representation from the absent modality with that of the existing modality, thereby mitigating the negative impact in the missing modality. Finally, the overall training objective for each client is formulated as follows:
$\mathcal{L}=\mathcal{L}_{S}+\alpha_{reg}\mathcal{L}_{CMPR}+\alpha_{con}\mathcal{L}_{CMPC}+\alpha_{align}\mathcal{L}_{CMA}$ (13)
@@
where $\alpha_{reg}$, $\alpha_{con}$ and $\alpha_{align}$ are hyper-parameters that can be tuned to control the importance of the $\mathcal{L}_{CMPR}$, $\mathcal{L}_{CMPC}$ and $\mathcal{L}_{CMA}$ terms, respectively. The overall training process for MFCPL is shown in Algorithm 1. The server broadcasts the complete prototypes to all clients in each communication round. In the local training phase, each client trains the learning model on private data using the overall training loss defined in Eq. (13).
Methods UCI-HAR (F1 score) Hateful Memes (AUC score) MELD (UAR score)
q = 0.5 q = 0.7 q = 0.8 q = 0.5 q = 0.7 q = 0.8 q = 0.5 q = 0.7 q = 0.8
FedAvg McMahan et al. (2017) 70.90 69.46 67.50 54.56 50.09 49.56 51.74 50.71 50.58
FedProx Li et al. (2020) 69.88 69.74 67.92 54.89 53.92 50.25 51.41 50.85 50.04
FedOpt Reddi et al. (2020) 74.79 66.17 56.45 54.62 51.61 49.58 50.96 49.28 48.27
MOON Li et al. (2021) 70.52 69.52 68.15 53.45 52.02 51.75 52.22 51.04 50.00
FedProto Tan et al. (2022b) 74.87 73.63 72.78 52.67 51.40 50.68 50.76 45.74 42.93
Ours (MFCPL) 77.44 75.61 75.19 56.30 55.69 55.04 53.73 52.24 51.92
@@ 4.1 Experimental Setup
  • UCI-HAR Anguita et al. (2013) dataset consists of smartphone sensory data from Accelerometer and Gyroscope sensors. This dataset involves observations from $30$ volunteers engaged in six distinct activities: walking, walking upstairs, walking downstairs, sitting, standing, and laying.
@@
  • Hateful Memes Kiela et al. (2020) dataset comprises $10,000$ samples of image and text pairs organized into binary classes. The primary objective is to identify and detect hateful speech within memes.
@@
  • MELD Poria et al. (2018) dataset is a multimodal conversational dataset for emotion recognition collected from Friends TV series, which contains about $13,000$ utterances from $1,433$ dialogues. Following the setting from Feng et al. (2023), we keep $4$ emotion labels with most samples which are neutral, sadness, happiness and anger.
@@
Following Feng et al. (2023), we initialize $50$, $86$, and $105$ clients with the participation rates of $0.25$, $0.1$, and $0.1$ per communication round for Hateful Memes, MELD, and UCI-HAR datasets, respectively. To simulate the non-IID scenario, we apply the Dirichlet distribution Ferguson (1973) with $\beta=0.2$ to the Hateful Memes and UCI-HAR datasets. In contrast, the MELD dataset inherently exhibits a natural non-IID data distribution due to its organization by speakerIDs.
Model Architecture.

### v2

@@ 3.2 Multimodal Prototype Federated Learning
Previous research work on federated prototype learning (Tan et al., 2022b; Mu et al., 2023) has focused on scenarios in which clients possess unimodal data. In the context of MFL with $M$ modalities and missing modality scenarios, multiple types of clients arise, such as those missing one or more modalities and those with $M$ modalities available. Simply aggregating unimodal prototypes based on modality may introduce a bias toward the dominant modality and cause the performance degradation in FL under severely missing modalities. This prompts us to leverage prototypes from all modalities to develop a generalized model for different types of clients.
Complete Prototypes
@@
$\{r_{\Omega_{i}}=g_{1}(e_{\Omega_{i}})\}\in\mathbb{R}^{d}.$ (2)
where ${\Omega_{i}}\subseteq\Omega=\{1,2,...,M\}$ denotes the subset of modalities available to a given client type. We define the $k^{th}$ class prototype from different types of clients as:
$p^{k}_{{\Omega_{i}}}=\frac{1}{|S^{k}_{{\Omega_{i}}}|}\sum_{i\in S^{k}_{{\Omega_{i}}}}r_{{\Omega_{i}}},$ (3)
where $S^{k}_{{\Omega_{i}}}$ represents the subset of $D_{i}$ belonging to $k^{th}$ class of different client types of $\Omega_{i}$. As complementary to unimodal prototypes, we introduce complete prototypes by averaging local prototypes from different types of clients in the shared representation space $\mathbb{R}^{d}$. Specifically, we formulate the complete prototypes on the server as follows:
$\begin{split}P^{k}&=\frac{1}{N}\sum_{i\in N}\{p^{k}_{{\Omega_{i}}}|\Omega_{i}\subseteq\Omega\}\in\mathbb{R}^{d}\\ P&=[P^{1},P^{2},\dots,P^{K}],\end{split}$ (4)
where $P^{k}$ denotes the complete prototypes belonging to class $k\in K$. Specifically, our complete prototypes guide the local training process with the calibration at two levels: modality-specific and modality-shared representations.
@@
3:        $\theta^{i}_{t},p_{{\Omega_{i}}}\leftarrow\textbf{LocalUpdate}(\theta_{t},P)$
@@
9:     for each batch $\in D_{i}=\{D_{i}=\{{x}^{1}_{i},{x}^{2}_{i},...,x^{M}_{i},{y}_{i}\}$ do
10:        $z_{m}=f_{m}(x^{m})$
11:        $e=h_{\text{self-attention}}([z_{1},z_{2},...z_{M}]),r_{\Omega_{i}}=g_{1}(e_{\Omega_{i}})$
12:        $z^{\prime}_{m_{i}}=g_{2}(z_{m_{i}})$
13:        $\mathcal{L}_{CMPR}\leftarrow(r_{\Omega_{i}},P)$ in Eq. (5)
14:        $\mathcal{L}_{CMPC}\leftarrow(z^{\prime}_{m_{i}},P)$ in Eq. (8)
15:        $\mathcal{L}_{CMA}\leftarrow(z^{\prime}_{m_{i}},\forall m\in\{1,2,...,M\})$ in Eq. (9)
@@
20:  $p_{\Omega_{i}}=[p^{1}_{\Omega_{i}},p^{2}_{\Omega_{i}},\dots,p^{K}_{\Omega_{i}}]$
21:  return $\theta^{i}_{t},p_{\Omega_{i}}$
3.3 Multimodal Federated Cross Prototype Learning
Cross-Modal Prototypes Regularization
Similar to other traditional FL methods, MFL encounters the challenge of data heterogeneity, and further complicated by the issue of missing modalities. In response to these challenges, we introduce a regularization loss that narrows the gap between local and global models, simultaneously providing diverse modality knowledge for clients with missing modalities at the modality-shared representation level. Drawing inspiration from FedProto (Tan et al., 2022b), we incorporate a regularization loss between local projected multimodal representations $r_{\Omega_{i}}$ and complete prototypes $P^{k}$ within the same semantic class $k$. However, our approach differs from FedProto, as our regularization loss serves as an extension to address the challenges posed by missing modalities in the context of multimodal federated learning. To implement this regularization, we utilize $\ell_{2}$ distance and introduce the Cross-Modal Prototypes Regularization (CMPR) as follows:
$\mathcal{L}_{CMPR}=\sum_{k}\|r^{k}_{\Omega_{i}}-P^{k}\|_{2}^{2},$ (5)
where $r^{k}_{\Omega_{i}}$ is the local projected multimodal representations of $k$ semantic class. By conducting calibration on the modality-shared representation with complete prototypes, we alleviate the impact of missing modalities from the multimodal perspective. Additionally, the regularization reduces the divergence between local models and global model, thereby mitigating data heterogeneity. Furthermore, we maintain the coherence between the modality-shared representation and complete prototypes in the shared space through the projection head.
Cross-Modal Prototypes Contrastive
@@ 3.3 Multimodal Federated Cross Prototype Learning
We envision that modality-specific representations in clients with missing modalities can capture rich and diverse modality knowledge from the complete prototype. The modality-specific features of local data instances should align closely with their respective complete prototypes within the same semantic class while exhibiting dissimilarity to complete prototypes associated with different semantic classes. For data instances $D_{i}=\{{x}^{1}_{i},{x}^{2}_{i},...,x^{M}_{i},{y}_{i}\}$, we encode each modality $m$ using its respective encoder $f_{m}$ to acquire the modality-specific feature vectors as follows:
$z_{m_{i}}=f_{m}({x}^{m}_{i})\in\mathbb{R}^{n},\forall m\in\{1,2,...,M\}.$ (6)
To facilitate alignment between the modality-specific features and complete prototypes, we employ the projection head $g_{2}$ to project the modality-specific feature vectors of each modality into the shared representation space:
$z^{\prime}_{m_{i}}=g_{2}(z_{m_{i}})\in\mathbb{R}^{d},\forall m\in\{1,2,...,M\}.$ (7)
@@
Let $p$ be the corresponding complete prototype $p\in P$, $C^{k}=P-P^{k}$ represents the prototypes with different semantic class from the local instances. Subsequently, the Cross-Modal Prototypes Contrastive (CMPR) loss where all modalities $M$ are available is expressed as follows:
$\mathcal{L}_{CMPC}=\sum_{m=1}^{M}-\log\frac{\sum_{p\in P^{k}}\exp(s(z^{\prime}_{m_{i}},p)/\tau)}{\sum_{p\in P}\exp(s(z^{\prime}_{m_{i}},p)/\tau)},$ (8)
@@
where $\mathrm{s}(u,v)=u^{\top}v/\lVert u\rVert\lVert v\rVert$ denotes the cosine similarity between the projected modality-specific feature and complete prototypes, and $\tau$ is the contrastive hyper-parameter. In cases where clients lack data for modality $m$, the corresponding contrastive loss for that modality $m$ will be excluded. The goal of Eq. (8) is to encourage the modality-specific representations from clients with missing modalities to closely align with the corresponding complete prototypes in the shared embedding space, thereby enhancing the generalization of these clients and mitigate the negative impact of missing modality from the unimodal perspective. Through this calibration, we utilize the projection head to facilitate alignment between features of the complete prototype and modality-specific representations. Additionally, we ensure the preservation of relevance between modality-specific representations and the complete prototypes in the shared space. The effectiveness of the projection head is further discussed in subsection 4.5.3.
(a) UCI-HAR
(b) Hateful Memes
(c) MELD
Figure 4: The data distribution of training data for each client with non-IID characteristic. The color bar denotes number of data samples. Each rectangle defines the number of samples in each class of each client.
Table 1: Client distribution on UCI-HAR, Hateful Memes and MELD datasets under various missing rates. I,T,A,V represent Image, Text, Audio, Video modalities, respectively.
Dataset Missing Rate Client Distribution
UCI-HAR 50% Acc: 35, Gyro: 40, Acc-Gyro: 30
70% Acc: 43, Gyro: 55, Acc-Gyro: 7
80% Acc: 54, Gyro: 48, Acc-Gyro: 3
100% Acc: 55, Gyro: 50
Hateful Memes 50% I: 20, T: 16, I-T: 14
70% I: 21, T: 26, I-T: 3
80% I: 30, T: 19, I-T: 1
100% I: 29, T: 21
MELD 50% A: 14, T: 13, V: 13, A-T: 14, A-V: 10, V-T: 10, A-T-V: 12
70% A: 26, T: 22, V: 22, A-T: 4, 6: 12, V-T: 5, A-T-V: 1
80% A: 28, T: 40, V: 23, A-T: 2, A-V: 3, V-T: 1
100% A: 24, T: 32, V: 30
Cross-Modal Alignment
@@
While the complete prototypes contribute diverse modality knowledge to address data heterogeneity and missing modalities at modality-specific and modality-shared representation levels, simply imputing missing data with zeros for the absent modality can introduce noise. We introduce an alignment loss between modality-specific representations from different modalities to alleviate negative effect arising from zero-filling of missing modalities. By aligning features within the shared representation space, the CMA loss facilitates effective complementarity between features from different modalities. We apply the alignment loss to the projected modality-specific representations. To achieve this goal, we compute the alignment loss using the $\ell_{2}$ distance between the projected representations of all available modality pairs. For $M$ modalities, the Cross-Modal Alignment loss is formulated as:
$\mathcal{L}_{CMA}=\sum_{(m_{1},m_{2})\in\mathcal{P}_{M}}\sum_{i\in D_{i}}\|z^{\prime}_{m_{1i}}-z^{\prime}_{m_{2i}}\|_{2}^{2},$ (9)
@@
where $\mathcal{P}_{M}$ denotes the set of all pairs of modalities among available $M$ modalities, and $z^{\prime}_{m_{1i}}=g_{2}(z_{m_{1i}})$, $z^{\prime}_{m_{2i}}=g_{2}(z_{m_{2i}})\in\mathbb{R}^{d}$ represent the projected modality-specific representations for modalities $m_{1}$ and $m_{2}$, respectively. Our objective is to align the representation from the absent modality with that of the existing modality, thereby mitigating the negative impact in the missing modality. Moreover, since the original representation spaces of each modality is different from each other, alignment in the shared space enables the coherence across modality-specific representation. This approach has shown effectiveness in improving performance in multimodal learning (Pham et al., 2019; Dong et al., 2023). Our experiments demonstrate that this alignment, as tailored in our design, significantly enhances the method’s effectiveness and illustrates its adaptability across various loss functions. Finally, the overall training objective for each client is formulated as follows:
$\mathcal{L}=\mathcal{L}_{S}+\alpha_{reg}\mathcal{L}_{CMPR}+\alpha_{con}\mathcal{L}_{CMPC}+\alpha_{align}\mathcal{L}_{CMA},$ (10)
@@
where $\alpha_{reg}$, $\alpha_{con}$ and $\alpha_{align}$ are hyper-parameters that control the importance of the $\mathcal{L}_{CMPR}$, $\mathcal{L}_{CMPC}$ and $\mathcal{L}_{CMA}$ terms, respectively. The overall training process for MFCPL is shown in Algorithm 1. The server broadcasts the complete prototypes to all clients in each communication round [^2] Iterative steps in federated learning where clients update local models and send them to server for aggregation. Each round represents one cycle of this process.. In the local training phase, each client trains the learning model on private data using the overall training loss defined in Eq. (10).
@@ 4.1.1 Datasets
    UCI-HAR (Anguita et al., 2013) dataset consists of smartphone sensory data from Accelerometer and Gyroscope sensors. This dataset involves observations from $30$ volunteers engaged in six distinct activities: walking, walking upstairs, walking downstairs, sitting, standing, and laying.
  • 2.
@@
    Hateful Memes (Kiela et al., 2020) dataset comprises $10,000$ samples of image and text pairs organized into binary classes. The primary objective is to identify and detect hateful speech within memes.
  • 3.
@@
    MELD (Poria et al., 2018) dataset is a multimodal conversational dataset for emotion recognition collected from Friends TV series, which contains about $13,000$ utterances from $1,433$ dialogues. In this experiment, we utilize audio, visual and text modalities. Following the setting from (Feng et al., 2023), we keep $4$ emotion labels with most samples which are neutral, sadness, happiness and anger.
@@
Following (Feng et al., 2023), we initialize $50$, $86$, and $105$ clients with the participation rates of $0.25$, $0.1$, and $0.1$ per communication round for Hateful Memes, MELD, and UCI-HAR datasets, respectively. We illustrate the data partition of three datasets in Figure 4. To simulate the non-IID scenario, we apply the Dirichlet distribution (Ferguson, 1973) with $\beta=0.2$ to the Hateful Memes and UCI-HAR datasets. In contrast, the MELD dataset inherently exhibits a natural non-IID data distribution due to its organization by speakerIDs. We perform a random split for the UCI-HAR (Anguita et al., 2013) dataset with $70\%$ allocated for training data and $30\%$ for test data. Within the training set, $20\%$ is further set aside for the validation set. In the case of the Hateful Memes dataset, we adhere to the setup outlined in the original work of this dataset (Kiela et al., 2020). Specifically, we construct a validation set and test set from $5\%$ and $10\%$ of total data, respectively, and use the remaining portion for training data. As for the MELD dataset, we adopt the train-test-validation split provided in the original work of this dataset (Poria et al., 2018). We utilize the validation procedure in model evaluation to select the optimal model and mitigate the risk of overfitting on the training set.
Methods UCI-HAR (F1 score) Hateful Memes (AUC score) MELD (UAR score)
$q=0.5$ $q=0.7$ $q=0.8$ $q=1.0$ $q=0.5$ $q=0.7$ $q=0.8$ $q=1.0$ $q=0.5$ $q=0.7$ $q=0.8$ $q=1.0$
FedAvg (McMahan et al., 2017) 70.90 69.46 67.50 66.85 54.56 50.09 49.56 47.66 51.22 50.62 50.36 47.46
FedProx (Li et al., 2020) 69.88 69.74 67.92 67.23 54.89 53.92 50.25 48.82 52.45 51.02 50.55 49.13
FedOpt (Reddi et al., 2020) 74.79 66.17 56.45 55.83 54.62 51.61 49.58 47.29 51.58 49.27 47.54 46.61
MOON (Li et al., 2021) 70.52 69.52 68.15 67.05 53.45 52.02 51.75 49.36 51.88 51.10 50.99 49.32
FedProto (Tan et al., 2022b) 74.87 73.63 72.78 70.99 52.67 51.40 50.68 47.15 50.66 47.51 46.94 36.80
FedProc (Mu et al., 2023) 70.14 69.37 68.77 65.43 52.75 50.49 50.23 48.81 51.92 51.83 50.48 49.80
FPL (Huang et al., 2023) 75.16 74.28 73.46 71.58 54.97 53.79 52.94 50.79 52.11 50.09 49.49 48.76
FedIoT (Zhao et al., 2022) 71.57 70.58 68.62 65.28 53.98 52.72 51.31 49.60 51.81 49.48 49.61 46.96
FedMSplit (Chen and Zhang, 2022) 74.76 73.32 72.77 71.65 55.57 53.61 52.65 51.80 52.97 51.65 50.57 48.29
Ours (MFCPL) 77.44 75.61 75.19 73.93 56.30 55.69 55.04 53.88 54.71 53.62 52.18 50.26
Table 3: Performance comparison of our MFCPL with SOTA methods on UCI-HAR, Hateful Memes and MELD datasets under various zero-filling data rates u with $q=0.5$. The best results are marked in bold.
Methods UCI-HAR (F1 score) Hateful Memes (AUC score) MELD (UAR score)
$u=0.2$ $u=0.5$ $u=0.7$ $u=0.2$ $u=0.5$ $u=0.7$ $u=0.2$ $u=0.5$ $u=0.7$
FedAvg (McMahan et al., 2017) 74.71 73.55 72.64 56.34 55.32 54.77 53.63 52.91 52.37
FedProx (Li et al., 2020) 74.34 73.35 71.97 57.67 56.78 55.47 53.82 53.18 52.88
FedOpt (Reddi et al., 2020) 78.45 77.08 75.86 57.52 56.91 55.21 53.90 52.69 51.79
MOON (Li et al., 2021) 73.45 72.57 73.27 57.03 56.52 54.50 53.55 53.02 52.86
FedProto (Tan et al., 2022b) 76.33 76.02 75.83 54.98 53.81 54.23 54.19 53.40 52.68
FedProc (Mu et al., 2023) 74.91 72.80 72.36 54.17 53.29 53.11 53.06 52.82 52.04
FPL (Huang et al., 2023) 76.92 76.17 75.23 55.59 55.26 54.61 54.38 53.45 52.83
FedIoT (Zhao et al., 2022) 75.82 74.13 72.03 56.27 55.16 54.19 53.87 52.64 51.89
FedMSplit (Chen and Zhang, 2022) 76.46 76.03 74.98 57.32 56.44 55.78 54.05 53.27 53.02
Ours (MFCPL) 78.84 77.87 78.03 58.35 57.49 56.83 55.36 55.06 54.81
4.1.2 Model Architecture